if the length of a recectangle is five less than three times its width and the perimeter is 105 inches, what is its width? length?
1
Expert's answer
2019-09-26T09:12:43-0400

Solution:

Suppose the length of the rectangle is : L

and Width of the Rectangle is : W


As per the the given question

L = 3W "-" 5 ---------- (1) // Length is five less than three times its width

Given Perimeter = 105 inches

2( L + W ) = 105 ---------- (2)

Put the value of L from equation (1) to equation (2)


2( 3W "-" 5 + W)= 105

6W "-" 10 + 2W = 105

8W "-" 10 = 105

8W = 115

W = 14.375


Put the value of W in the equation (1)

L = 3W "-" 5

L = 3"*"14.375 "-" 5

L = 43.125 "-" 5

L = 38.125


Answer : W = 14.375 inches

L = 38.125 inches